OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = (x – 4)(2x – 1)(x + 7)?

OpenStudy (anonymous):

(x – 4)(2x – 1)(x + 7) =0 x-4=0 or 2x-1=0 or x+7=0 x=4 or x==1/2 or x=-7

OpenStudy (anonymous):

By definition, a zero is a number you plug into the function so that f(x)=0. They are also called roots. So if f(r)=0, then r is a root and a zero of the function.